OpenStudy (anonymous):

Find the Domain and range and graph. \[y=1/2\sqrt{x}+3\]

OpenStudy (anonymous):

is it \[\frac{1}{2\sqrt x}||\frac{1}{2}\sqrt x\]

OpenStudy (anonymous):

im not sure what you're asking

OpenStudy (anonymous):

domain is easy just look at the denominator and figure out what makes it equal 0 you can't divide by 0 to find the range just convert it to its inverse (if it has an inverse) and use that to determine the range

OpenStudy (anonymous):

you cant take the square root of a negative number either so that is another limitation on your domain

OpenStudy (anonymous):

since this only has positive numbers it is clear that it passes the horizontal line test so you can take the inverse to find its range

OpenStudy (anonymous):

@cuddlepony so i set y to 0?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the denominator of the function cannot equal 0 because oyu cannot divide by zero thus any x value that renders the denominator 0 is not in the domain. Furthermore, you have a radical, you can't take the square root of a negative number thus it limits your domain further.

OpenStudy (anonymous):

to find the range just find the inverse of your function (if possible) and determine its domain as that will provide you with your range

OpenStudy (anonymous):

I don't understand how to get the domain then, it can't equal 2?

OpenStudy (anonymous):

 is this your function?

OpenStudy (anonymous):

nope.